# Avoided-crossings, degeneracies and Berry phases in the spectrum of quantum noise through analytic Bloch-Messiah decomposition

### v2

@@ Abstract
The Bloch-Messiah decomposition (BMD) is a fundamental tool in quantum optics, enabling the analysis and tailoring of multimode Gaussian states by decomposing linear optical transformations into passive interferometers and single-mode squeezers. Its extension to frequency-dependent matrix-valued functions, recently introduced as the “analytic Bloch-Messiah decomposition” (ABMD), provides the most general approach for characterizing the driven-dissipative dynamics of quantum optical systems governed by quadratic Hamiltonians. In this work, we present a detailed study of the ABMD, focusing on the typical behavior of parameter-dependent singular values and of their corresponding singular vectors. In particular, we analyze the hitherto unexplored occurrence of avoided and genuine crossings in the spectrum of quantum noise, the latter being manifested by nontrivial topological Berry phases of the singular vectors. We demonstrate that avoided crossings arise naturally when a single parameter is varied, leading to hypersensitivity of the singular vectors and suggesting the presence of genuine crossings in nearby systems. We highlight the possibility of programming the spectral response of photonic systems through the deliberate design of avoided crossings. As a notable example, we show that such control can be exploited to generate broad, flat-band squeezing spectra – a desirable feature for enhancing degaussification protocols. This study provides new insights into the structure of multimode quantum correlations and offers a theoretical framework for experimental exploitation of complex quantum optical systems.
I Introduction
Bloch-Messiah decomposition (BMD) allows to break down “any optical circuit with linear input-output relations into a linear multiport interferometer followed by a unique set of single-mode squeezers and then another multiport interferometer” and establishes that “squeezing is an irreducible resource which remains invariant under transformations by linear optical elements” [1]. It is, therefore, a pivotal technique enabling the understanding and manipulation of multimode Gaussian states as well as the development of practical applications for quantum optics and the processing of quantum information with continuous variables (CV). For example it plays a central role in the generation and exploitation of continuous-variables cluster states in optical frequency combs [2, 3, 4, 5] or in multimode spatial systems [6, 7, 8]. As well, it is used for engineering complex multimode quantum states [9, 10, 11, 12]. In simple passage nonlinear optics, BMD has been employed for the modal analysis of pulsed parametric-down conversion (PDC) and frequency conversion (FC) [13, 14], of temporal imaging [15], of coherent mode-selective photon subtraction [16, 17, 18] and amplification with four-wave mixing (FWM) [19]. In general, the Schmidt decomposition of the joint spectral amplitude in PDC and FC can be seen as a special case of BMD [20, 21].
When the nonlinear dynamics of a system of $n$ boson modes occurs in a cavity, it is still possible to apply BMD to the system’s transfer function evaluated at a given working frequency. This approach, referred to as Schmidt mode decomposition in the work by Gonzalez-Arciniegas et al. [22], has been employed to study the quantum noise of light generated in $\chi^{(3)}$-based optical parametric oscillators (OPOs) [22, 23] or the classical noise in pulsed $\chi^{(2)}$-based OPOs [24]. However, this analysis is not complete as it was first recognized by Jiang et al. in [25], who studied pulsed $\chi^{(2)}$-based OPOs where the pump is characterized by a non-trivial carrier-envelope offset (CEO). Under the simplifying assumption of a compensated-dispersion cavity configuration, it was observed that a set of supermodes depending on a continuous CEO parameter $\theta$ must be considered to fully describe the time/frequency properties of the generated non-classical light.
@@ I Introduction
All these systems actually belong to a larger and more general class of problems that are characterized by quantum Langevin equations generated by quadratic Hamiltonians. Gouzien et al. [26] developed a general approach based on the “analytic Bloch-Messiah decomposition” (ABMD). In this framework, the system’s transfer function is decomposed in terms of frequency-dependent supermodes, referred to as “morphing supermodes” and physically intepreted as “interferometers with memory effect” (IME) [27], with the singular values representing the associated frequency-dependent squeezing levels. ABMD has been instrumental in characterizing “hidden squeezing” [28] and identifying an optimal measurement strategy for its detection based on IMEs. It is remarkable that an approach similar to ABMD, referred to as “Frequency-dependent principal component analysis”, was recently developed for characterizing neuronal dynamics in brain activity [29].
@@
One of the merits of this work is to provide an in-depth description of the fundamental properties of the ABMD, and to highlight its role in the analysis of mode dynamics in quantum optical systems. To this end, we review the generic behavior of singular values and singular vectors in one-parameter families of matrices, and specialize this analysis to the case of interest: the ABMD. In doing so, we elucidate that the typical phenomenon encountered is that of avoided crossings of singular values, where the singular values appear to be about to coalesce but suddenly veer away from each other. Avoided crossings are typically accompanied by sharp variations in the corresponding singular vectors, a behaviour that makes the unitary factors of the ABMD highly sensitive to small perturbations. In contrast, true intersections of singular values – referred to as degeneracies in this work, as diabolical points in the physics literature, and as conical intersections in the chemistry literature – are generally not expected to occur unless the number of parameters is increased to match the codimension [^1] Roughly speaking, the codimension of a phenomenon refers to the number of parameters that must be varied for the phenomenon to be generically observed. For a rigorous treatment of this concept, we refer the reader to [62]. Throughout this work, dimension and codimension will always be considered with respect to the real field. of the degeneracy. Another contribution of this work is to clarify that the term ABMD actually encompasses a family of decompositions, all sharing the same singular values but generally differing in their unitary factors. We then focus our attention on a specific decomposition, which we name joint-minimum-variation BMD – or ABMD, depending on the smoothness of the matrix-valued function. We present an algorithm for computing it numerically (caution: smoothly computing a BMD/ABMD is not just collecting BMDs at separate parameter values, see Remark 3), and describe how it can be used to reveal avoided crossings and locate nearby degeneracies by exploiting certain topological properties of the unitary factors. All these concepts are illustrated through concrete examples of realistic physical systems in which such behaviors emerge, thereby demonstrating their relevance in experimental settings.
We note that much of our analysis builds upon earlier work. A fundamental reference is the classical monograph by Kato [31], which provides a comprehensive treatment of the spectral theory of parameter-dependent operators. Particularly relevant to our study are also the contributions by Bunse-Gerstner et al. [32] on the numerical computation of the analytic singular value decomposition, as well as a series of works by Dieci and co-authors [33, 34, 35, 36, 37, 38, 39, 40], in which various decompositions (notably, eigendecompositions and singular value decompositions) of one-parameter and multi-parameter families of matrices are systematically investigated, with special attention to their numerical construction and their use in detecting spectral degeneracies.
@@
A very interesting physical consequence of the occurrence of avoided crossings is the possibility of exploiting them - particularly the avoided crossing - to spectrally engineer the quantum noise properties of the associated singular vectors. As pointed out by Asavanant et al. in [41], the quality of de-Gaussification of continuous-variable (CV) squeezed states depends on the flatness of the squeezing spectrum, which has so far required the use of narrowband filters. However, this necessarily reduces the amount of nonclassical correlations available for the preparation of non-Gaussian states. We show that avoided crossings can be exploited to tailor quantum states featuring a broad, nearly flat-band squeezing spectrum.
II The playground: linear quantum Langevin equations and their decoupling
@@
[…]
@@
$S$ is “$\omega$-symplectic” which stands for a smooth (possibly analytic) matrix-valued function that is conjugate symplectic for all $\omega$, see Definition 1. This follows from the fact that $\mathcal{M}$ is an Hamiltonian matrix, i.e. ${(\Omega\,\mathcal{M})}^{\mathrm{T}}=\Omega\,\mathcal{M}$, and $\Gamma$ is a skew-Hamiltonian, i.e. ${(\Omega\,\Gamma)}^{\mathrm{T}}=-\Omega\,\Gamma$.
@@
A matrix-valued function of $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{p})\in\mathbb{R}^{p}$ is said to be “$\bm{\lambda}$-symplectic” if it is smooth (possibly analytic) and conjugate symplectic for all values of $\bm{\lambda}$.
@@
This “symplectic diagonalization” exists only for $\mathcal{M}$ symmetric [44, 45, 46], which is equivalent to having $G=0$.
@@
III Smooth decompositions
III.1 Static SVD
@@
The degree of non-uniqueness of the SVD depends on the number of distinct singular values. If the singular values of $A\in\mathbb{C}^{n\times n}$ are all distinct, its singular vectors are unique (only) up to an arbitrary phase factor. That is, if $A=U\Sigma V^{\dagger}$ is an SVD of $A$ with $\sigma_{1}>\sigma_{2}>\ldots>\sigma_{n}$, then any other SVD must be of the form $A={\widetilde{U}}\Sigma{\widetilde{V}}^{\dagger}$, where $\widetilde{U}=U\Theta$ and $\widetilde{V}=V\Theta$, with $\Theta=\operatorname{diag}\big{(}e^{i\theta_{1}},e^{i\theta_{2}},\ldots,e^{i\theta_{n}}\big{)}$ for some $\theta_{j}\in\mathbb{R}$, $j=1,\ldots,n$. However, if $A\in\mathbb{C}^{n\times n}$ has fewer than $n$ distinct singular values, then the unitary factors of its SVD gain additional degrees of freedom. For instance, suppose that $A=U\Sigma V^{\dagger}$ is an SVD of $A$ with $\sigma_{1}=\sigma_{2}>\sigma_{3}>\ldots>\sigma_{n}$. Then, $A={\widetilde{U}}\Sigma{\widetilde{V}}^{\dagger}$ is an SVD of $A$ if and only if $\widetilde{U}=U\begin{bmatrix}W&0\\ 0&\Theta\end{bmatrix}$ and $\widetilde{V}=V\begin{bmatrix}W&0\\ 0&\Theta\end{bmatrix}$, where $W$ is a $2\times 2$ complex unitary matrix and $\Theta=\operatorname{diag}\big{(}e^{i\theta_{3}},e^{i\theta_{4}},\ldots,e^{i\theta_{n}}\big{)}$ with $\theta_{j}\in\mathbb{R}$, $j=3,\ldots,n$. If the number of distinct singular values drops below $n-1$, the freedom in choosing the unitary factors of the SVD increases further.
III.2 Dynamic SVD
@@ III.3 Codimension of degeneracies
It is known that $\mathcal{S}_{n}$, as a subset of $\mathbb{C}^{n\times n}$, has real dimension $32n^{2}-$ 3 (see [33]), where $2n^{2}$ is just the real dimension of $\mathbb{C}^{n\times n}$. Hence, the real codimension of $\mathcal{S}_{n}$ is $3$ footnotemark: . To see why, it is insightful to look at the $2\times 2$ case, where $\mathcal{S}_{2}$ is just the set of positive scalar multiples of $2\times 2$ unitary matrices:
$\mathcal{S}_{2}=\{\sigma W,\text{ with }\sigma\geq 0\text{ and }W\in\mathbb{C}^{2\times 2}\text{ unitary }\},$
@@
The codimension of the degenerate set $\mathcal{S}_{n}$ may decrease due to additional symmetries in the matrices. For example, the codimension is 2 for real symmetric matrices (see [35]) and complex symmetric matrices (see [39]). Another instance of reduced codimension will be encountered later, in Section IV.
III.4 Static and dynamic Bloch-Messiah decomposition
From this point onward, we focus on conjugate symplectic matrices (see Definition (1)). For consistency with the notation introduced in Section II, we denote the singular values of a ${2n\times 2n}$ conjugate symplectic matrix by $d_{1},\ldots,d_{n},d_{1}^{-1},\ldots,d_{n}^{-1}$, where $d_{1}\geq\ldots\geq d_{n}\geq 1$, and denote by $D$ the diagonal matrix of its singular values:
@@ III.4 Static and dynamic Bloch-Messiah decomposition
where each block is ${n\times n}$, e.g. see [50, p. 14]; it is also easy to see that a conjugate symplectic real diagonal matrix $D\in\mathbb{R}^{2n\times 2n}$ must have the form $D=\operatorname{diag}\left(D_{1},D_{1}^{-1}\right)$.
@@
0.  $S\in\mathbb{C}^{2n\times 2n}$ conjugate symplectic with distinct singular values.
0.  $U,V\in\mathbb{C}^{2n\times 2n}$ and $D_{1}\in\mathbb{R}^{n\times n}$ as in Theorem 1.
@@
Let $S(\omega)$, $\omega\in\mathbb{R}$, be a $\omega$-symplectic matrix function (see Property 2) having distinct singular values for all $\omega$, and consider a smooth Bloch-Messiah decomposition of $S$ as in Theorem 2. We will call (16) a joint-minimum-variation BMD over the interval $[0,\tau]$ if the pair $(U,V)$ minimizes the quantity
[…]
@@
As we have already seen, the inherent non-uniqueness of the unitary factors in the BMD allows for (infinitely) many different ways of constructing a smooth BMD for a given $\omega$-symplectic matrix-valued function. We remark that the recipe used by Gouzien et al. (see supplemental notes in  [26], page 5), which amounts to imposing the condition $\operatorname{diag}(V(\omega)^{\dagger}\dot{V}(\omega))=0$, corresponds  [^2] This follows from Theorem 2.6 in [37]. to minimizing only the variation of the left unitary factor $V$. Consequently, the resulting ABMD differs from the joint-minimum-variation ABMD, in which the variations of both unitary factors contribute to the minimized quantity. Both recipes are equally valid if one is solely interested in the spectrum of the singular values. However, as we will illustrate in Section IV, the joint-minimum-variation BMD is the mandatory choice for detecting degeneracies, especially when their codimension may be lower than anticipated.
@@
over the interval $[0,2\pi]$. Note that each $\alpha_{j}$ can be chosen to be a continuous function of $\theta$ (see [37]). If $\alpha_{j}(\pi)\neq\alpha_{j}(0)$ for some $j$, then $S$ must be degenerate at some point ${\bm{\lambda}}_{0}$ inside the sphere $\mathbb{S}$, and the degeneracy must involve the singular value $\sigma_{j}$.
@@
In Section IV, Theorem 3 will be used as a topological tool to detect parameter values where certain symplectic matrix functions are degenerate. In order to do so, we will need to compute the joint-minimum-variation BMD of some matrix functions. Below, we outline the algorithm we have used to accomplish this task. The algorithm is an adaptation of the one proposed in [38] to our context.

### v3

# Avoided crossings, degeneracies and Berry phases in the spectrum of quantum noise of driven-dissipative bosonic systems

@@ Abstract
Avoided crossings are fundamental phenomena in quantum mechanics and photonics that originate from the interaction between coupled energy levels and have been extensively studied in linear dispersive dynamics. Their manifestation in open, driven-dissipative systems, however, where nonlinear dynamics of quantum fluctuations come into play, remains largely unexplored. In this work, we analyze the hitherto unexplored occurrence of avoided and genuine crossings in the spectrum of quantum noise. We demonstrate that avoided crossings arise naturally when a single parameter is varied, leading to hypersensitivity of the associated singular vectors and suggesting the presence of genuine crossings (diabolical points) in nearby systems. We show that these spectral features can be deliberately designed, highlighting the possibility of programming the quantum noise response of photonic systems. As a notable example, such control can be exploited to generate broad, flat-band squeezing spectra – a desirable feature for enhancing degaussification protocols. Our analysis is based on a detailed study of the Analytic Bloch-Messiah Decomposition (ABMD), which we use to characterize the parameter-dependent behavior of singular values and their corresponding vectors. This study provides new insights into the structure of multimode quantum correlations and offers a theoretical framework for the experimental exploitation of complex quantum optical systems.
1 Introduction
Avoided crossings are ubiquitous phenomena in physics, particularly in quantum mechanics and condensed matter, where they provide a hallmark of mechanisms such as quantum interference and energy-level repulsion [44, 33, 55]. They have been extensively studied in contexts such as molecular spectroscopy [24, 51], quantum dots [32, 43], superconductors [42, 45], and semiconductor nanostructures [46, 47]. In photonics, avoided crossings emerge from the interaction between distinct optical modes, leading to the characteristic splitting and mutual repulsion of their resonant frequencies. These effects occur in a wide variety of optical systems—including waveguides [29, 54], photonic crystals [36], microresonators [23], and coupled-cavity architectures [4, 41]—and can be precisely engineered through the control of coupling strengths and system parameters. This control enables functionalities ranging from mode conversion and enhanced light–matter coupling to spectral filtering and dispersion engineering [30, 17, 34].
Such avoided crossings are features of the linear energy spectrum and arise from number-conserving quadratic interactions (e.g., beam-splitter-like couplings). In other words, they originate from mode-hopping terms associated with linear dispersive dynamics. In nonlinear photonic platforms—such as Kerr microresonators—these linear avoided crossings reshape the dispersion landscape and can be exploited to enable or tailor phenomena such as dispersive-wave emission and soliton formation [49, 50, 53]. However, in those cases the avoided crossing remains a property of the underlying linear mode structure, rather than a dynamical feature of the nonlinear quantum evolution itself.
@@ 1 Introduction
It is far from obvious that such spectral features persist once the full driven–dissipative dynamics—including nonlinear or parametric processes—is taken into account in the presence of quantum noise. In this regime, after linearization around the classical steady state, quantum fluctuations are governed by quantum Langevin equations generated by an effective quadratic Hamiltonian that includes pair-generation terms. The evolution of quantum fluctuations is governed by smooth, frequency-dependent conjugate-symplectic transformations—which we term $\omega$-symplectic transformations for brevity. The appropriate tool for their analysis is the Analytic Bloch–Messiah Decomposition (ABMD), which provides a smooth factorization of these transformations. This framework was introduced in [20], developing a general approach where the system’s transfer function is decomposed in terms of frequency-dependent supermodes. These modes—referred to as morphing supermodes and physically interpreted as interferometers with memory effect (IME) [16]—have singular values that represent the corresponding frequency-dependent squeezing levels. The ABMD framework has been instrumental in revealing hidden squeezing [21] and in identifying an optimal measurement strategy for its detection based on IMEs. Remarkably, a conceptually similar method, termed frequency-dependent principal component analysis, has recently been proposed to characterize neuronal dynamics in brain activity [6].
1.1 Contributions
In this work, we provide an in-depth analysis of the dynamics in driven–dissipative bosonic systems and show the occurrence of avoided crossings in the spectrum of quantum noise, where the squeezing spectra approach each other and then abruptly veer apart.
Avoided crossings are typically accompanied by sharp variations in the corresponding singular vectors, a behaviour that makes them highly sensitive to small perturbations. In contrast, true intersections of singular values – referred to as degeneracies in this work, as diabolical points in the physics literature, and as conical intersections in the chemistry literature – are generally not expected to occur unless the number of parameters is increased to match the codimension [^1] Roughly speaking, the codimension of a phenomenon refers to the number of parameters that must be varied for the phenomenon to be generically observed. For a rigorous treatment of this concept, we refer the reader to [25]. Throughout this work, dimension and codimension will always be considered with respect to the field of real numbers. of the degeneracy.
A proper understanding of these features requires a formalism capable of tracking both singular values and their associated unitary factors in a smooth and physically consistent way. This role is naturally fulfilled by the Analytic Bloch–Messiah Decomposition (ABMD). We provide a detailed description of its fundamental properties and highlight its relevance in the analysis of mode dynamics.
@@ 1.1 Contributions
Another contribution of this work is to clarify that the term ABMD actually encompasses a family of decompositions, all sharing the same singular values but generally differing in their unitary factors. We then focus our attention on a specific decomposition, which we name joint-minimum-variation BMD – or ABMD, depending on the smoothness of the matrix-valued function. We present an algorithm for computing it numerically (caution: smoothly computing a BMD/ABMD is not just collecting BMDs at separate parameter values, see Remark 3), and describe how it can be used to reveal avoided crossings and locate nearby degeneracies by exploiting certain topological properties of the unitary factors. All these concepts are illustrated through concrete examples of realistic physical systems in which such behaviors emerge, thereby demonstrating their relevance in experimental settings.
@@
An interesting physical implication of these avoided crossings is the possibility of exploiting them to spectrally engineer the quantum-noise properties of the associated singular vectors. As pointed out by Asavanant et al. in [1], the quality of de-Gaussification of continuous-variable (CV) squeezed states depends on the flatness of the squeezing spectrum. Achieving this flatness has so far required narrowband filters, which, however, necessarily reduce the amount of nonclassical correlations available for the preparation of non-Gaussian states. We show that avoided crossings can be exploited to tailor quantum states featuring a broad, nearly flat-band squeezing spectrum.
2 The playground: linear quantum Langevin equations and their decoupling
@@
$\hat{H}=\hbar\sum_{j,\ell}G_{j,\ell}\hat{a}_{j}^{\dagger}\hat{a}_{\ell}+\frac{\hbar}{2}\sum_{j,\ell}[F_{j,\ell}\hat{a}_{j}^{\dagger}\hat{a}^{\dagger}_{\ell}+\text{h.c.}],$ (1)
@@
$S$ is $\omega$-symplectic” which stands for a smooth (possibly analytic) matrix-valued function that is conjugate symplectic for all $\omega$, see Definition 1. This follows from the fact that $\mathcal{M}$ is an Hamiltonian matrix, i.e. ${(\Omega\,\mathcal{M})}^{\mathrm{T}}=\Omega\,\mathcal{M}$, and $\Gamma$ is a skew-Hamiltonian, i.e. ${(\Omega\,\Gamma)}^{\mathrm{T}}=-\Omega\,\Gamma$.
@@
A matrix-valued function of $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{p})\in\mathbb{R}^{p}$ is said to be $\bm{\lambda}$-symplectic” if it is smooth (possibly analytic) and conjugate symplectic for all values of $\bm{\lambda}$.
@@
This “symplectic diagonalization" exists only for $\mathcal{M}$ symmetric [8, 18, 22], which is equivalent to having $G=0$.
@@
3 Smooth decompositions
3.1 Static SVD
@@
The degree of non-uniqueness of the SVD depends on the number of distinct singular values. If the singular values of $A\in\mathbb{C}^{n\times n}$ are all distinct, its singular vectors are unique (only) up to an arbitrary phase factor. That is, if $A=U\Sigma V^{\dagger}$ is an SVD of $A$ with $\sigma_{1}>\sigma_{2}>\ldots>\sigma_{n}$, then any other SVD must be of the form $A={\widetilde{U}}\Sigma{\widetilde{V}}^{\dagger}$, where $\widetilde{U}=U\Theta$ and $\widetilde{V}=V\Theta$, with $\Theta=\operatorname{diag}\big(e^{i\theta_{1}},e^{i\theta_{2}},\ldots,e^{i\theta_{n}}\big)$ for some $\theta_{j}\in\mathbb{R}$, $j=1,\ldots,n$. However, if $A\in\mathbb{C}^{n\times n}$ has fewer than $n$ distinct singular values, then the unitary factors of its SVD gain additional degrees of freedom. For instance, suppose that $A=U\Sigma V^{\dagger}$ is an SVD of $A$ with $\sigma_{1}=\sigma_{2}>\sigma_{3}>\ldots>\sigma_{n}$. Then, $A={\widetilde{U}}\Sigma{\widetilde{V}}^{\dagger}$ is an SVD of $A$ if and only if $\widetilde{U}=U\begin{bmatrix}W&0\\ 0&\Theta\end{bmatrix}$ and $\widetilde{V}=V\begin{bmatrix}W&0\\ 0&\Theta\end{bmatrix}$, where $W$ is a $2\times 2$ complex unitary matrix and $\Theta=\operatorname{diag}\big(e^{i\theta_{3}},e^{i\theta_{4}},\ldots,e^{i\theta_{n}}\big)$ with $\theta_{j}\in\mathbb{R}$, $j=3,\ldots,n$. If the number of distinct singular values drops below n1n-1, the freedom in choosing the unitary factors of the SVD increases further.
3.2 Dynamic SVD
@@ 3.3 Codimension of degeneracies
It is known that $\mathcal{S}_{n}$, as a subset of $\mathbb{C}^{n\times n}$, has real dimension $2n^{2}-3$ (see [9]), where $2n^{2}$ is just the real dimension of $\mathbb{C}^{n\times n}$. Hence, the real codimension of $\mathcal{S}_{n}$ is 33 footnotemark: . To see why, it is insightful to look at the $2\times 2$ case, where $\mathcal{S}_{2}$ is just the set of positive scalar multiples of $2\times 2$ unitary matrices:
[…]
@@
The codimension of the degenerate set $\mathcal{S}_{n}$ may decrease due to additional symmetries in the matrices. For example, the codimension is 2 for real symmetric matrices (see [13]) and complex symmetric matrices (see [11]). Another instance of reduced codimension will be encountered later, in Section 4.
3.4 Static and dynamic Bloch-Messiah decomposition
From this point onward, we focus on conjugate symplectic matrices (see Definition (1)). For consistency with the notation introduced in Section 2, we denote the singular values of a $2n\times 2n$} conjugate symplectic matrix by $d_{1},\ldots,d_{n},d_{1}^{-1},\ldots,d_{n}^{-1}$, where $d_{1}\geq\ldots\geq d_{n}\geq 1$, and denote by $D$ the diagonal matrix of its singular values:
@@ 3.4 Static and dynamic Bloch-Messiah decomposition
where each block is $n\times n$}, e.g. see [35, p. 14]; it is also easy to see that a conjugate symplectic real diagonal matrix $D\in\mathbb{R}^{2n\times 2n}$ must have the form $D=\operatorname{diag}\left(D_{1},D_{1}^{-1}\right)$.
@@
[…]
@@
Let $S(\omega)$, $\omega\in\mathbb{R}$, be a $\omega$-symplectic matrix function (see Property 2) having distinct singular values for all $\omega$, and consider a smooth Bloch-Messiah decomposition of $S$ as in Theorem 2. We will call (16) a joint-minimum-variation BMD over the interval $0,\tau]$ if the pair $(U,V)$ minimizes the quantity
$\int_{0}^{\tau}\sqrt{\norm{\dot{U}(\omega)}^{2}_{\mathrm{F}}+\norm{\dot{V}(\omega)}^{2}_{\mathrm{F}}}\,d\omega\,$, (18)
@@
As we have already seen, the inherent non-uniqueness of the unitary factors in the BMD allows for (infinitely) many different ways of constructing a smooth BMD for a given $\omega$-symplectic matrix-valued function. We remark that the recipe used by Gouzien et al. (see supplemental notes in [20], page 5), which amounts to imposing the condition $\operatorname{diag}(V(\omega)^{\dagger}\dot{V}(\omega))=0$, corresponds  [^2] This follows from Theorem 2.6 in [14]. to minimizing only the variation of the left unitary factor $V$. Consequently, the resulting ABMD differs from the joint-minimum-variation ABMD, in which the variations of both unitary factors contribute to the minimized quantity. Both recipes are equally valid if one is solely interested in the spectrum of the singular values. However, as we will illustrate in Section 4, the joint-minimum-variation BMD is the mandatory choice for detecting degeneracies, especially when their codimension may be lower than anticipated.
@@
over the interval $0,2\pi]$. Note that each $\alpha_{j}$ can be chosen to be a continuous function of $\theta$ (see [14]). If $\alpha_{j}(\pi)\neq\alpha_{j}(0)$ for some $j$, then $S$ must be degenerate at some point ${\bm{\lambda}}_{0}$ inside the sphere $\mathbb{S}$, and the degeneracy must involve the singular value $\sigma_{j}$.
@@
In Section 4, Theorem 3 will be used as a topological tool to detect parameter values where certain symplectic matrix functions are degenerate. In order to do so, we will need to compute the joint-minimum-variation BMD of some matrix functions. Below, we outline the algorithm we have used to accomplish this task. The algorithm is an adaptation of the one proposed in [10] to our context.